OpenStudy (anonymous):

A bag contains 4 red, 3 green, and 5 yellow marbles. Three marbles are drawn, one at a time without replacement. Determine the probability that the order in which they are selected is a) yellow, red, green b) yellow, green, green c) yellow, yellow, red

OpenStudy (anonymous):

i realy need help

OpenStudy (kropot72):

a) First draw: P(yellow) = 5/12 Second draw: P(red) = 4/11 Third draw: P(green) = 3/10 The result of each draw is independent of the other results. Therefore the overall probability of yellow, red, green is the product of the individual probabilities: \[P(yellow,\ red,\ green)=\frac{5}{12}\times \frac{4}{11}\times \frac{3}{10}\]

OpenStudy (anonymous):

thnx for helping me

OpenStudy (kropot72):

You're welcome :)